How to solve the following problem?
d^2 - d - 30 = 0

Answer by algebrahouse.com(1659)

dē - d - 30 = 0
(d - 6)(d + 5) = 0 {factored into two binomials}
d - 6 = 0 or d + 5 = 0 {set each factor equal to 0}
d = 6 or d = 5 {solved each equation for d}

For more help from me, visit: www.algebrahouse.com